# Truth or Twist? Optimal Model Selection for Reliable Label Flipping Evaluation in LLM-based Counterfactuals

### v1

@@ Appendix F Automatic Evaluation: A Sanity Check
Qwen2.5-7B Qwen2.5-7B $62.22$ $85.80$ $66.66$
DeepSeek-R1-Distill-Qwen-7B $61.60$ $93.33$ $80.00$
Qwen2.5-72B $62.22$ $84.44$ $71.11$
BERT $31.11$ $86.66$ $60.00$
RoBERTa $37.77$ $85.40$ $68.88$
Phi4-14B $66.66$ $93.33$ $66.66$
Mistral-Large $64.44$ $88.88$ $60.00$
User Study $80.00$ $62.22$ $100.00$
Qwen2.5-14B Qwen2.5-14B $57.77$ $93.33$ $84.44$
@@
Llama3-8B Llama3-8B $60.20$ $70.20$ $75.55$
DeepSeek-R1-Distill-Llama-8B $58.20$ $73.80$ $71.11$
Llama3-70B $53.00$ $72.20$ $82.22$
BERT $39.00$ $62.00$ $71.11$
RoBERTa $38.00$ $69.80$ $62.22$
Phi4-14B $60.00$ $70.20$ $68.88$
Mistral-Large $52.00$ $67.80$ $62.22$
User Study $75.56$ $46.67$ $100.00$
Llama3-70B Llama3-70B $62.60$ $85.60$ $84.44$
@@
Qwen2.5-7B Qwen2.5-7B $31.11$ $77.77$ $71.11$
DeepSeek-R1-Distill-Qwen-7B $48.88$ $80.00$ $55.55$
Qwen2.5-72B $37.77$ $80.00$ $60.00$
BERT $24.44$ $68.88$ $67.07$
RoBERTa $32.00$ $71.11$ $60.00$
Phi4-14B $35.55$ $79.24$ $57.77$
Mistral-Large $53.00$ $73.33$ $50.15$
User Study $77.78$ $53.33$ $100.00$
Qwen2.5-14B Qwen2.5-14B $26.66$ $88.88$ $62.22$
@@
Llama3-8B Llama3-8B $42.22$ $80.00$ $64.44$
DeepSeek-R1-Distill-Llama-8B $40.00$ $82.22$ $55.55$
Llama3-70B $33.33$ $73.33$ $73.33$
BERT $15.55$ $71.11$ $77.77$
RoBERTa $17.77$ $73.33$ $75.55$
Phi4-14B $40.00$ $80.00$ $80.00$
Mistral-Large $31.11$ $75.55$ $80.00$
User Study $86.67$ $73.33$ $100.00$
Llama3-70B Llama3-70B $35.55$ $91.11$ $62.22$
@@
Table 4: Label flip rate (in %) for counterfactual examples generated by Qwen2.5-{7B,14B,32B} and Llama3-{8B,70B}, evaluated across judge models exhibiting same, distilled, same family, and independent w/ and w/o fine-tuning relationships, and user study on 45 selected counterfactuals (§4.5) each from AG News, SST2 and SNLI datasets. Orange-highlighted values indicate that the judge model with the given relationship aligns closely with human annotators, while purple-highlighted values indicate the opposite.
In our user study, we randomly select 45 counterfactuals from each dataset, generated individually by five counterfactual generator models $LLM_{\mathcal{G}}$4.5). These samples are then provided to human annotators to evaluate the effectiveness of the counterfactual, known as the simulatability test Wiegreffe et al. (2021); Feldhus et al. (2023). To perform a sanity check and validate the representativeness of these subsets, we conduct an additional automatic evaluation on the subset of input data from which the 45 counterfactuals were generated.
@@
Appendix G Counterfactual Data Augmentation
@@
G.1 Evaluation on the original test set
@@
Qwen2.5-7B-Instruct Qwen2.5-7B-Instruct 0.831 0.875 0.645 0.262 0.565 0.440
DeepSeek-R1-Distill-Qwen-7B 0.839 0.823 0.614 0.291 0.579 0.258
Qwen2.5-72B-Instruct 0.842 0.858 0.612 0.274 0.546 0.302
BERT 0.881 0.837 0.567 0.277 0.561 0.320
RoBERTa 0.872 0.858 0.519 0.254 0.554 0.249
Phi4-14B 0.857 0.853 0.627 0.259 0.549 0.387
Mistral-Large-Instruct 0.838 0.858 0.653 0.247 0.556 0.333
Qwen2.5-14B-Instruct Qwen2.5-14B-Instruct 0.845 0.817 0.637 0.281 0.578 0.364
DeepSeek-R1-Distill-Qwen-14B 0.837 0.822 0.621 0.274 0.546 0.258
Qwen2.5-72B-Instruct 0.838 0.811 0.656 0.291 0.561 0.280
BERT 0.857 0.857 0.600 0.274 0.558 0.316
RoBERTa 0.869 0.791 0.632 0.284 0.550 0.262
Phi4-14B 0.825 0.825 0.629 0.284 0.576 0.338
Mistral-Large-Instruct 0.803 0.788 0.642 0.257 0.542 0.222
Qwen2.5-32B-Instruct Qwen2.5-32B-Instruct 0.813 0.836 0.646 0.277 0.547 0.347
DeepSeek-R1-Distill-Qwen-32B 0.842 0.836 0.648 0.286 0.533 0.351
Qwen2.5-72B-Instruct 0.774 0.831 0.646 0.264 0.532 0.316
BERT 0.788 0.852 0.594 0.247 0.540 0.351
RoBERTa 0.866 0.853 0.644 0.274 0.569 0.267
Phi4-14B 0.856 0.827 0.653 0.301 0.540 0.396
Mistral-Large-Instruct 0.831 0.780 0.631 0.269 0.523 0.298
@@
Llama3-70B-Instruct Meta-Llama-3-70B-Instruct 0.853 0.803 0.606 0.267 0.595 0.347
@@
Table 5: Downstream task performance (measured in terms of accuracy) is evaluated on test set and the set of counterfactuals (out-of-distribution instances) after applying counterfactual data augmentation on a BERT model ($LLM_{\mathcal{C}}$). The training data consist of original examples from the target dataset, along with counterfactual examples generated by Qwen2.5-7B,14B,32B and Llama3-8B,70B using FIZLE. The counterfactual labels are provided by different judge models exhibiting various relationships: same model, distilled, same family, and independent models with fine-tuning or without fine-tuning, across the AG News, SST2, and SNLI datasets.
G.2 Evaluation on the counterfactual set
In comparison to Section §G.1, where $LLM_{\mathcal{B}}$ and $LLM_{\mathcal{C}}$ are evaluated on the test set of each dataset, we further evaluate the fine-tuned BERT model $LLM_{\mathcal{C}}$ for CDA on the set of 45 selected counterfactuals, whose labels are obtained through the user study (§4.5). Note that the 45 selected counterfactuals and their corresponding original input texts are excluded from the training data for the model.
@@ G.2 Evaluation on the counterfactual set
We calculate the Spearman correlation between the rankings of the generator and judge models in Table 1 and Table 5 (counterfactual (CF) set). The results show a moderate correlation of 0.41 on the AG News dataset, indicating that the relationships might impact the performance of the CAD. Specifically, a better relationship may lead to higher accuracy on the CF test set. In contrast, a weak correlation is observed on other datasets.

### v2

@@ G.3 A Sanity Check
Llama3-8B
Llama3-8B $60.20$ $70.20$ $75.55$
DeepSeek-R1-Distill-Llama-8B $58.20$ $73.80$ $71.11$
Llama3-70B $53.00$ $72.20$ $82.22$
BERT $39.00$ $62.00$ $71.11$
RoBERTa $38.00$ $69.80$ $62.22$
Phi4-14B $60.00$ $70.20$ $68.88$
Mistral-Large $52.00$ $67.80$ $62.22$
User Study $75.56$ $46.67$ $100.00$
Qwen2.5-14B
Qwen2.5-14B $57.77$ $93.33$ $84.44$
@@
Llama3-70B
Llama3-70B $62.60$ $85.60$ $84.44$
@@
Llama3-8B
Llama3-8B $42.22$ $80.00$ $64.44$
DeepSeek-R1-Distill-Llama-8B $40.00$ $82.22$ $55.55$
Llama3-70B $33.33$ $73.33$ $73.33$
BERT $15.55$ $71.11$ $77.77$
RoBERTa $17.77$ $73.33$ $75.55$
Phi4-14B $40.00$ $80.00$ $80.00$
Mistral-Large $31.11$ $75.55$ $80.00$
User Study $86.67$ $73.33$ $100.00$
Qwen2.5-14B
Qwen2.5-14B $26.66$ $88.88$ $62.22$
@@
Llama3-70B
Llama3-70B $35.55$ $91.11$ $62.22$
@@
Table 5: Label flip rate (in %) for counterfactual examples generated by Qwen2.5-{14B,32B} and Llama3-{8B,70B}, evaluated across judge models exhibiting same, distilled, same family, and independent w/ and w/o fine-tuning relationships, and user study on 45 selected counterfactuals (§4.4) each from AG News, SST2 and SNLI datasets. Orange-highlighted values indicate that the judge model with the given relationship aligns closely with human annotators, while purple-highlighted values indicate the opposite.
In our user study, we randomly select 45 counterfactuals from each dataset, generated individually by five counterfactual generator models $LLM_{\mathcal{G}}$4.4). These samples are then provided to human annotators to evaluate the effectiveness of the counterfactual, known as the simulatability test Wiegreffe et al. (2021); Feldhus et al. (2023). To perform a sanity check and validate the representativeness of these subsets, we conduct an additional automatic evaluation on the subset of input data from which the 45 counterfactuals were generated.
@@
Appendix H Counterfactual Data Augmentation
@@
H.1 Evaluation on the original test set
@@
Qwen2.5-14B-Instruct
Qwen2.5-14B-Instruct 0.845 0.817 0.637 0.281 0.578 0.364
DeepSeek-R1-Distill-Qwen-14B 0.837 0.822 0.621 0.274 0.546 0.258
Qwen2.5-72B-Instruct 0.838 0.811 0.656 0.291 0.561 0.280
BERT 0.857 0.857 0.600 0.274 0.558 0.316
RoBERTa 0.869 0.791 0.632 0.284 0.550 0.262
Phi4-14B 0.825 0.825 0.629 0.284 0.576 0.338
Mistral-Large-Instruct 0.803 0.788 0.642 0.257 0.542 0.222
Qwen2.5-32B-Instruct
Qwen2.5-32B-Instruct 0.813 0.836 0.646 0.277 0.547 0.347
DeepSeek-R1-Distill-Qwen-32B 0.842 0.836 0.648 0.286 0.533 0.351
Qwen2.5-72B-Instruct 0.774 0.831 0.646 0.264 0.532 0.316
BERT 0.788 0.852 0.594 0.247 0.540 0.351
RoBERTa 0.866 0.853 0.644 0.274 0.569 0.267
Phi4-14B 0.856 0.827 0.653 0.301 0.540 0.396
Mistral-Large-Instruct 0.831 0.780 0.631 0.269 0.523 0.298
Llama3-70B-Instruct
Meta-Llama-3-70B-Instruct 0.853 0.803 0.606 0.267 0.595 0.347
@@
Table 6: Downstream task performance (measured in terms of accuracy) is evaluated on test set and the set of counterfactuals (out-of-distribution instances) after applying counterfactual data augmentation on a BERT model ($LLM_{\mathcal{C}}$). The training data consist of original examples from the target dataset, along with counterfactual examples generated by Qwen2.5-{14B,32B} and Llama3-{8B,70B} using FIZLE. The counterfactual labels are provided by different judge models exhibiting various relationships: same model, distilled, same family, and independent models with fine-tuning or without fine-tuning, across the AG News, SST2, and SNLI datasets.
H.2 Evaluation on the counterfactual set
In comparison to Section §H.1, where $LLM_{\mathcal{B}}$ and $LLM_{\mathcal{C}}$ are evaluated on the test set of each dataset, we further evaluate the fine-tuned BERT model $LLM_{\mathcal{C}}$ for CDA on the set of 45 selected counterfactuals, whose labels are obtained through the user study (§4.4). Note that the 45 selected counterfactuals and their corresponding original input texts are excluded from the training data for the model.
@@ H.2 Evaluation on the counterfactual set
We calculate the Spearman correlation between the rankings of the generator and judge models in Table 3 and Table 6 (CF set). The results show a moderate correlation of 0.41 on the AG News dataset, indicating that the relationships might impact the performance of the CAD. Specifically, a better relationship may lead to higher accuracy on the CF test set. In contrast, a weak correlation is observed on other datasets.